You are saving for retirement. On April 1, you deposit $515 in your target-date fund. By May 15, your initial deposit is worth $530 and you deposit an additional $262 at that time. On August 15, your account shows a value of $826 and you make a further deposit of $271. On November 30, your account shows a balance of $1087. What is your time-weighted return (in percent) for the period April 1 - November 30?   Answer to two decimals. Carry intermediate calculations to six decimals.
Answer #1

Answer:

Holding Period Return (HPR) = (End Value - Initial Value + Cash Flow) / (Initial Value + Cash Flow)

HPR ( April 1 to May 15) = (530 - 515) / 515 = 2.912621%

HPR (May 15 to August 15) = (826 - (530 + 262))/ (530 +262) = 4.292929%

HPR (August 15 to November 30) = (1087 - (826 + 271))/ (826 + 271) = - 0.911577%

Time-weighted return (in percent) for the period April 1 - November 30

= ((1 + 2.912621%) * (1 + 4.292929%) * (1 - 0.911577%)) -1

= 6.352186%

= 6.35%

Time-weighted return (in percent) for the period April 1 - November 30 = 6.35%